# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

### v3

@@ Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge in the computational sciences. Traditional sampling methods like molecular dynamics (MD) or MCMC often struggle with the high-dimensional nature of molecular systems and the high energy barriers of transitions between stable states. While these transitions are rare events in simulation timescales, they often represent the most biologically significant processes, for example, the conformational change of an ion channel protein from its closed to open state, which controls cellular ion flow and is crucial for neural signaling. Such transitions in real systems may take milliseconds to seconds but could require months or years of continuous simulation to observe even once. We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation. By representing transition paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular dynamics force fields, our method enables the efficient discovery of physically realistic transition pathways without requiring expensive path sampling. We demonstrate our method’s effectiveness on two proteins, including an explicitly hydrated bovine pancreatic trypsin inhibitor (BPTI) system with over 8,300 atoms. Our approach reproduces the conformational change captured in a landmark millisecond-scale explicit-solvent MD simulation (Shaw et al., 2010), while achieving remarkable computational efficiency, requiring only force field evaluations compared to the approximately 412 billion evaluations in the original study. This represents a nearly six orders of magnitude reduction, allowing us to generate the transition pathway in just 15 minutes on a standard GPU rather than weeks on specialized hardware.
@@ 1 Introduction
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable transition paths to minimizers of specific action functionals in different limiting regimes. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths through a discrete series of molecular configurations, connected by artificial spring forces to ensure continuity or as splines. Although grounded in solid theoretical principles, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates. The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints. Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for many small-scale systems, it has not yet fully benefited from modern optimization approaches that could address these challenges.
@@
Concurrent advances in scientific machine learning and molecular simulation have created opportunities for novel approaches to this challenge. Physics-informed neural networks (PINNs) have revolutionized our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019). Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical forcefield models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives needed for path optimization.
@@
The last decade has seen numerous approaches emerge to tackle the transition path sampling problem. Traditional shooting and other sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, require initial MD-harvested trajectories to begin with or preexisting knowledge of the transition dynamics, and face inherent tradeoffs between rejection rates and path diversity. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample full transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024). Notably, Du et al. (2024) use a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble via Doob’s h-transform. This approach provides rich statistical information about transition mechanisms. Most closely related to our method, Ramakrishnan et al. (2025) employ a continuous neural parameterization of minimum-energy paths inspired by Nudged Elastic Band (NEB), incorporating specialized components for tangential forces, path velocity, and transition state energetics. Another promising concurrent approach by Raja et al. (2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a forcefield within a chain-of-states framework to generate plausible transition distributions. While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to large biomolecular systems with explicit solvent while maintaining fast training times.
@@
Here, we introduce an approach focused on efficient minimum-energy path discovery as a foundation for understanding molecular transition mechanisms. Our method represents transition paths as continuous, differentiable neural functions that map a progress coordinate to molecular configurations. This representation, inspired by recent advances in neural implicit representations (Mildenhall et al., 2020; Sitzmann et al., 2020), naturally incorporates physical constraints through a physics-informed neural network (PINN) framework. By directly targeting the minimum-energy path through energy minimization, we achieve remarkable computational efficiency, enabling applications to complex biomolecular systems in minutes on standard hardware while extracting the most physically relevant information about transition mechanisms. The resulting pathways can serve as valuable starting points for established transition path sampling methods (Dellago et al., 1998; Bolhuis et al., 2002) when exploration of the complete ensemble is desired.
@@ 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To our loss function, we lean on alternative formulations of the Onsager-Machlup action functional as derived by Vanden-Eijnden & Heymann (2008); Olender & Elber (1997) in the context of the string methods. We begin with a molecular system evolving under overdamped Langevin dynamics, a well-established model for molecular motion in the high-friction regime characteristic of biomolecular systems:
@@
To obtain the probability of an entire continuous trajectory, we take the product of these conditional probabilities over all time steps and then take the limit as . This yields the path probability:
The exponent term can be recognized as the Onsager-Machlup action functional, which is defined as:
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. The action penalizes paths that deviate from following the deterministic drift . To further simplify, we expand the integrand of the action functional:
Applying the Cauchy-Schwarz inequality to the kinetic and potential terms, we can establish an important relationship:
with equality holding if and only if is parallel (or antiparallel) to . When the path is reparameterized so that velocity consistently aligns with the gradient, the action is minimized, yielding the most likely transition pathway. For such pathways where is everywhere proportional to , we can transform the time integral into a purely geometric one. However, to maintain this parallel alignment, the path length itself must be subject to optimization, as a fixed-length path would generally not satisfy the parallelity condition. By defining the arc-length element along the spatial curve traced out by , we obtain:
This geometric reformulation connects directly to the concept of a minimum-energy path (MEP), which follows the potential energy landscape’s gradient while avoiding high-energy regions. The MEP represents the most probable transition mechanism in the overdamped limit, making it an ideal target for our optimization.
We represent the transition path as a continuous neural mapping for such that and , where represents the network parameters. To transform the geometric line integral into a practical loss function, we first discretize the path and apply Taylor expansion to relate gradient magnitudes to energy differences:
where and . Using the triangle inequality, we can establish that minimizing provides an upper bound on this action, which yields our neural network loss function:
A complete derivation with all mathematical details is provided in Appendix A.1. By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
@@ 2.1.1 Neural Representation of minimum-energy Paths
Drawing inspiration from the string method formalism, we propose representing the transition path using an implicit neural parameterization, similar to the mean parameterization of Du et al. (Du et al., 2024). While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively). Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms:
Here is a neural network with parameters , and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the network’s output. While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective. The blending function controls the network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
@@ 2.1.2 Path Optimization
Rather than alternating between evolution and reparameterization steps as in the classical string method, we minimize directly through stochastic gradient descent. For each batch of paths:
The implicit neural representation naturally handles reparameterization constraints through its continuous parameterization. Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary—the inherent continuity of , combined with smooth blending functions, ensures path continuity without additional constraints.
@@ 2.2.1 String Method
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Ramakrishnan et al., 2025) determines MEPs by evolving a curve in configuration space according to physical forces while maintaining proper parameterization. Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized by that connects these states. The MEP satisfies:
@@ 2. Reparameterization step to enforce equal spacing:
This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition . Unlike the neural approach presented in this paper, which handles parameterization implicitly, this method requires explicit reparameterization steps that can introduce computational inefficiencies.
@@ 2.2.2 Chain-of-States Methods
where and is a spring constant. The spring term enforces continuity of the path while allowing flexibility in the discretization. However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions. The neural representation described earlier avoids this limitation through its natural continuity properties.
At convergence, both formulations yield a discretized MEP satisfying at each image, the same condition that PINN-MEP targets through energy minimization, but with the advantage of continuous representation and gradient-based optimization.
@@ 3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix A.2, we compare and benchmark our method to the related approach outlined in (Du et al., 2024), and in Appendix B.1, we present comprehensive ablation studies analyzing the impact of different embedding strategies, interpolation methods, network architectures, and regularization approaches.
@@ 4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. Beyond its therapeutic applications, BPTI has become a fundamental model system for studying protein dynamics due to its moderate size (58 residues) and complex conformational changes involving disulfide bond rearrangements.
@@
To investigate BPTI’s conformational transitions, we analyze five key states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare insight into BPTI’s transition mechanisms. Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
@@
We generate a transition MEP connecting the initial and final BPTI states from the D.E. Shaw simulation. The intermediate snapshots from the D.E. Shaw simulation serve as validation points; they were not used as inputs but represent physically realized conformations from the whole trajectory. To analyze this path, we compute two collective variables matching those used in the original D.E. Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD of residues 4-54 after mean-centering. Remarkably, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive explicit-solvent MD simulation.
@@
We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation. We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations. To handle water molecules during transitions, we pair proximal water molecules between start and end states, then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network. This hybrid approach allows us to include explicit solvent effects while keeping computational costs low. The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
@@
The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996). Each minimum-energy path was computed in approximately 15 minutes on a single A6000 GPU, compared to the D.E. Shaw simulation, which required approximately 1.3 million node-hours on their specialized Anton supercomputer to generate the 1.03 millisecond trajectory capturing the same conformational changes. Appendix B.2.6 provides a detailed computational performance comparison.

### v4

@@ Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high energy barriers between stable states. These rare events often represent the most biologically significant processes, like ion channel opening that controls cellular signaling, yet may require months of continuous simulation to observe. We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation. By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling. We demonstrate our approach in two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms. Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study. Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in just 15 minutes on a standard GPU rather than weeks on specialized hardware.
@@ 1 Introduction
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths through a discrete series of molecular configurations, connected by artificial spring forces to ensure continuity or as splines. Although grounded in solid theoretical principles, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates. The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints. Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for many small-scale systems, it has not yet fully benefited from modern optimization approaches that could address these challenges.
@@
Concurrent advances in scientific machine learning and molecular simulation have created opportunities for novel approaches to this challenge. Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019). Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical forcefield models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives needed for path optimization.
@@
The last decades have seen numerous approaches emerge to tackle the transition path sampling problem. Traditional shooting and other sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, require initial MD-harvested trajectories to begin with or preexisting knowledge of the transition dynamics. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample full transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024). Notably, Du et al. (2024) use a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble via Doob’s h-transform. This approach provides rich statistical information about transition mechanisms. Most closely related to our method, Ramakrishnan et al. (2025) employ a continuous neural parameterization of minimum-energy paths inspired by Nudged Elastic Band (NEB), incorporating specialized components for tangential forces, path velocity, and transition state energetics. Another promising concurrent approach by Raja et al. (2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a forcefield within a chain-of-states framework to generate plausible transition distributions. While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to large biomolecular systems with explicit solvent while maintaining fast training times.
@@
Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy path finding problem using a continuous optimization framework. Our approach leverages the universal function approximation capabilities of neural networks to represent conformational transitions as differentiable mappings from a progress coordinate to molecular configurations. By embedding physical principles directly into the optimization objective through a physics-informed neural network (PINN) framework, we eliminate the need for explicit reparameterization steps that burden traditional methods. This design enables applications to complex biomolecular systems in minutes on standard hardware while extracting physically meaningful transition mechanisms. The resulting pathways can serve as valuable starting points for established transition path sampling methods (Dellago et al., 1998; Bolhuis et al., 2002) when exploration of the complete ensemble is desired.
@@ 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To derive our loss function, which when minimized yields a continuous curve representing the minimum-energy path connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods. However, we further simplify the approach after reaching the geometric formulation. We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
@@
Taking the product of these conditional probabilities over all time steps and then taking the limit as yields the path probability:
The exponent term can be recognized as the Onsager-Machlup action functional Onsager & Machlup (1953):
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. To further simplify, we expand the integrand of the action functional:
For paths with fixed endpoints, the cross-term integrates to a constant difference in potential energy:
Since and are fixed in our formulation, this term becomes a constant that doesn’t affect the optimization. We can therefore focus on minimizing the remaining terms. Applying the Cauchy-Schwarz inequality:
This upper bound becomes an equality when is parallel or antiparallel to , with the direction determined by whether the path is ascending or descending the energy landscape. Crucially, in our formulation, we allow the path to optimize its shape and effective length. This freedom enables the path to adjust to follow the energy gradient (or its opposite), making the above upper bound tight at the optimum. When we abandon fixed-time parameterization and allow the path to stretch as needed to align with energy gradients, we can transform the time integral into a purely geometric one by defining the arc-length element along the spatial curve :
[…]
where and .
For a path segment between adjacent points, applying Taylor’s theorem, we can express the energy difference as:
Where is the angle between the gradient and displacement vectors. For minimum-energy paths, the displacement aligns with or against the gradient direction () giving:
This gives us a surrogate objective with the same minimum:
Using the triangle inequality on the absolute energy differences, we can establish that minimizing the simple sum of energies along the path provides an upper bound on this action:
Since the latter terms are constants during optimization, this leads to our final loss function. To implement this in practice, we sample B points along a parametrized path and directly minimize the sum of energies at these points, yielding:
where are uniformly or randomly sampled points in . A complete derivation with all mathematical details is provided in Appendix A.1. By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
@@ 2.1.1 Neural Representation of Minimum-Energy Paths
We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al. (Du et al., 2024). While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively). Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms:
Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the network’s output. While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective. The blending function controls the network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
@@ 2.1.2 Path Optimization
We minimize directly through stochastic gradient descent. For each batch of points along the path:
Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary: the inherent continuity of , combined with smooth blending functions, ensures path continuity without additional constraints.
@@ 2.2.1 String Method
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Ramakrishnan et al., 2025) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization. Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized by that connects these states. The MEP satisfies:
@@ 2. Reparameterization step to enforce equal spacing:
This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition . This method requires explicit reparameterization steps that can introduce computational inefficiencies.
@@ 2.2.2 Chain-of-States Methods
where and is a spring constant. The spring term enforces the continuity of the path while allowing flexibility in the discretization. However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions. Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost. The neural representation avoids this limitation through its natural continuity properties.
@@ 3 Experiments
We evaluate our method on two molecular systems: the AIB9 peptide and Bovine Pancreatic Trypsin Inhibitor (BPTI). These systems represent different scales of complexity and serve to validate our method’s ability to discover physically meaningful transition paths. In Appendix A.2, we compare and benchmark our method with the related approach outlined in (Du et al., 2024), and in Appendix B.1, we present ablation studies analyzing the impact of different embedding strategies, interpolation methods, network architectures, and regularization approaches.
@@ 4 BPTI Conformational Change Pathway
Bovine Pancreatic Trypsin Inhibitor (BPTI) is a clinically significant protein used as an anticoagulant in medical procedures. BPTI is 58 residues long and has complex conformational changes involving disulfide bond rearrangements paired with backbone changes.
@@
To investigate BPTI’s conformational transitions, we use as reference structures five key states from a landmark molecular dynamics (MD) simulation performed by D.E. Shaw Research on their Anton supercomputer (Shaw et al., 2010). These publicly available snapshots provide rare benchmark structures for validating our approach to BPTI’s transition mechanisms. Full details of our BPTI system preparation, explicit solvent shell construction, and hybrid neural network architecture for handling water molecules are provided in Appendix B.2.
@@
We generate multiple transition MEPs connecting two BPTI states from the D.E. Shaw simulation. The intermediate snapshots from the D.E. Shaw simulation serve as validation points; they were not used as inputs but represent physically realized conformations from the whole trajectory. To analyze this path, we compute two collective variables matching those used in the original D.E. Shaw study: (1) the disulfide torsion angle between Cys14 and Cys38, calculated using the dihedral angle formed by CB14–SG14–SG38–CB38 atoms, and (2) the backbone RMSD of residues 4-54 after mean-centering. Remarkably, our optimized path passes through all intermediate snapshots, indicating that it successfully discovers the same transition mechanism as the computationally intensive explicit-solvent MD simulation.
@@
We matched their simulation conditions by incorporating explicit water molecules in our model to ensure a meaningful comparison with the Anton simulation. We generated water shells extending 14Å from any protein atom for both the initial and final BPTI conformations. To handle water molecules during transitions, we pair proximal water molecules between start and end states, then model their movement using simple parameterized splines, rather than directly incorporating all water coordinates into the neural network. This hybrid neural-spline approach allows us to include explicit solvent effects while keeping computational costs low. The specific details of our water shell construction and matching algorithm are described in Appendix B.2.
@@
The system was modeled using the AMBER99sb-ILDN force field (Lindorff-Larsen et al., 2010) with the TIP3P water model (Jorgensen et al., 1983; Neria et al., 1996). Each minimum-energy path was computed in approximately 15 minutes on a single A6000 GPU, compared to the D.E. Shaw simulation, which required approximately 1.3 million node-hours on their specialized Anton supercomputer to generate the 1.03 millisecond trajectory capturing the same conformational changes. While our approach does not provide the same statistical ensemble information as equilibrium MD, it efficiently captures the key transition pathway. It’s worth noting that even the reference millisecond-scale simulation may not have fully achieved equilibrium sampling of the conformational space. Appendix B.2.7 provides a detailed computational performance comparison.